OpenStudy (anonymous):

what is the distance between (13,7) (-9,4)

OpenStudy (tad1):

Use the distance formula D = [sqrt[{y2-y1)^2 + (x2-x1)^2}]

OpenStudy (anonymous):

distance = root {(x2-x1)^2 + (y2-y1)^2}

OpenStudy (anonymous):

root {(x2-x1)^2 + (y2-y1)^2} \[ =\sqrt{[13-(-9)]^2+[7-4]}=\sqrt{(22^{2})+(3^{2})}=\sqrt{493}=22.2\]

OpenStudy (anonymous):

got it???